# How Do Developers Interact with AI? An Exploratory Study on Modeling Developer Programming Behavior

### 28 Mar 2026 v1

@@ 2.6. Annotation Tool and Procedure
To collect data for the S-IASE model, we developed an annotation tool so participants can divide video recordings into multiple time intervals and annotate each interval based on the model.
Hence, we used the implementation of a tool for annotating programming videos developed by Mozannar et al.’s (Mozannar et al., 2024) as our starting point. We designed the graphical user interface (GUI) and also developed features that are essential for our study purpose. For instance, we improved the user experience of the annotation tool by displaying the options with dropdown menus following the pilot study.
Figure 2 depicts our annotation tool GUI. Once the task execution stage is finished, we asked participants to stop recording their screens. We then provided a demo on the annotation tool. After that, participants annotated their programming session to ensure their recollection of the task remained fresh. Unlike the original tool (Mozannar et al., 2024), ours allows participants to define their own time intervals by setting start and end timestamps.
@@
In the annotation step, we encouraged participants to annotate approximately 20 intervals (each representing a different time range) describing their development states. This guideline was based on the average number of intervals observed during our pilot study. Although we used 20 intervals as a reference for participants, they could freely create as many intervals as they felt necessary. For each time interval, participants needed to annotate all four dimensions of our S-IASE model (described in Section 4).
In the initial annotation tool, the predefined categories contained 10 intentions, 14 actions, 8 supporting tools, and 7 emotions labels. They could also add custom labels within each category if they found predefined options insufficient. After data collection, two researchers first reviewed a random subset of 10 screen videos and their corresponding annotated labels to assess the completeness and consistency. After validation, the same researchers conducted a structured merging (Zimmermann, 2016) process to consolidate similar custom categories and refine categories based on participant responses. One researcher reviewed and clustered all custom labels, mapping them onto existing categories or creating new categories as necessary (e.g., adding the intention to set up project environment and the action search). Another researcher independently validated these mappings, and any disagreements (i.e., representing 5% of all categories) were resolved through discussions until consensus was achieved. As shown in Table 3, the newly emerged labels are marked with a (+), indicating that they were added during this merging stage.
@@ 2.7. Pilot Study
Before experimenting on a larger group of participants, we performed a pilot study with eight participants for two purposes: (1) design and refine our developer-AI model derived from our literature review (which was S-IAS at the time, no E), and (2) refine details in our designed user study steps to ensure clarity for participants. The pilot study became pivotal for answering RQ1 because we discovered through the pilots that, in addition to intention and action, emotion emerged as a paramount dimension participants repeatedly alluded to. For example, “I am so frustrated that ChatGPT is granting permissions to all the repos via the REST API as opposed to having a more privacy-centric view in its answers.” It became clear to us from the pilots that participants demonstrated a fluctuating level of emotions while using AI tools, which are important to consider if we want to improve and strengthen prior characterizations of developer-AI interaction during programming. This finding is in line with prior work that has studied the impact of emotion on developer productivity (Madampe et al., 2025; Cheng et al., 2023; Graziotin et al., 2015; Girardi et al., 2020).
As a result, we made a major refinement to our model, adding emotion as a dimension.
This resulted in four dimensions that formed the foundation for our S-IASE model. Additionally, piloting helped identify annotation categories for each of these four dimensions, for example, the intention To Clarify Requirements/Ambiguities emerged when participants expressed uncertainty regarding task instructions. Actions such as Logging and tools like Database Tools were added to better capture task-specific activities observed during the piloting sessions.
Piloting also allowed us to refine our task instructions to improve clarity and feasibility. For example, the 25-minute task duration was determined based on the performance of pilot participants, who demonstrated that the core requirements of both PyT and JavaT could be addressed within this time-boxed window. For JavaT, we allowed participants to choose between Maven or standard package imports to establish a JDBC connection, reducing setup complexity. We also improved the annotation tool UI by making it into a single page application with all the annotation categories for a given time interval in a scrollable panel. To prevent data contamination, we confirm pilot participants were not included in the main study results.
@@ 2.8. Interview
As the last stage of our user study, we invited participants to our follow-up interview. Before the interview, the authors proposed a set of general interview questions from multiple perspectives.
For constructing our interview scripts, we mainly focused on what they thought, why they behaved in certain ways, how they felt during the experiments, and what they have learned. Therefore, the interview scripts were constructed in three layers. First, we designed general questions on developer–AI interaction inspired by prior studies (e.g., (Mozannar et al., 2024; Barke et al., 2023; Li et al., 2024; Liang et al., 2024; Choudhuri et al., 2024)) and adopted Bloom’s taxonomy (Bloom et al., 1956; Anderson and Krathwohl, 2001)
@@
to probe what participants felt they had learned during the task. Second, we focused on cognitive biases and cognitive load under time pressure, using responses from validated instruments such as the NASA-TLX (Hart, 2006) and self-efficacy scales as anchors for follow-up questions (e.g., “You rated mental demand as 5, what was happening in that moment?”). Third, we included questions on our annotation tool usability, asking participants to evaluate our annotation tool and whether it accurately captured their intentions, actions, and emotions. Then, for each interview, we carefully watched the participant’s uploaded video and read their responses history to tailor participant-specific follow-up “why” questions, such as “Why did you switch from using an AI tool to searching Stack Overflow at this point?”
@@
Among the 76 participants, 14 (18%) volunteered to be interviewed. Each interview took around 30 minutes with two authors and one participant. Our interviews followed a semi-structured format where we used our prepared interview questions as a general guideline, but allowed flexibility to explore topics that emerged in an interview.
After obtaining permission from our interviewees, interview audio was recorded and transcribed by Otter.ai (50). To ensure that no errors were made in the transcription, one author manually verified the transcript for each interview.
@@
The same author who conducted each interview and transcription afterwards performed basic data analysis from socio-technical grounded theory (STGT) (Hoda, 2024). We chose STGT for data analysis due to its grounding in the socio-technical context. Part of the reason STGT was developed was due to “the core SE practice of programming is not strictly technical, rather, it is socio-technical, involving intensive human-human collaboration and coordination and human-technology interactions.” (Hoda, 2021) With the addition of AI-assistance to programming, the subject matter we studied is fittingly in the socio-technical domain. In contrast to traditional grounded theory methods, which are dedicated to theory development and do not support limited applications for data analysis, STGT can support data analysis to generate rich descriptive themes and categories. These themes and categories helped supplement our proposed developer-AI model and findings about observed developer behavior. Furthermore, we adopted a reflexive, researcher-led approach to conduct socio-technical grounded theory (STGT) from an interpretivist perspective (Hoda, 2024). The first author served as the primary analytic instrument, conducting open coding, constant comparison, and iterative memoing across all interview transcripts. This process resulted in a range of themes. For this paper, we present themes most aligned with our research focus, such as “Trust but Verify”. Member checking with a subset of 6 interviewees helped confirm the resonance and validity of the presented interpretations.
@@ 3. Task Completion Results
Completed Corrected Completed Corrected Completed Corrected Completed Corrected
@@
Completed Corrected Completed Corrected Completed Corrected Completed Corrected Completed Corrected
@@
Table 1 details how our participants performed on each of the subtasks for PyT and JavaT. We distinguished between Completed (the participant performed the required actions and finished the specific subtask) and Corrected (the outcome of the subtask was verified to be functionally accurate). A participant achieved “All Subtasks Corrected” if they completed and correctly solved all subtasks within 25 minutes. Among 39 participants for PyT, only 8 of them were successful in solving all subtasks, while 2 out of 37 for JavaT. We found that all 10 participants who correctly solved all their assigned subtasks are from AI-assisted groups. We found that 42% (32/76) of all participants successfully reached the final subtask of their assigned activity (56% for PyT and 27% for JavaT). However, the overall rate for full functional correctness remained at 13% (10/76). This suggests that while the tasks were possible to navigate within the time limit, the complexity of API and database interactions presented a substantial challenge to functional accuracy. The lower success rate in the Java task, compared to the Python task, further indicates that environment setup and multi-step database operations imposed a heavier correctness burden on developers. Moreover, we found that the AI group produced a higher completion rate than the non-AI group for each subtask. The results also indicate that while AI markedly boosted both completion and correctness, the complexity of the tasks presented a substantial challenge within the 25-minute limit.
On average, participants in AI-assisted groups adopted more than one AI assistant (1.37) during the experiment. ChatGPT was by far the most popular, used by 38 participants, with an additional 13 participants relying on the paid version. 12 participants used GitHub Copilot as AI assistants in IDE. This distribution indicated multiple assistants’ strategies for adopting AI for task completion.
@@ 4.1. Outcome
Our participants were able to add labels to the model as needed. After data collection and structured merging, we had 12 intentions, 16 actions, 8 supporting tools, and 7 emotions. Newly emerged labels are marked with (+) to indicate they were added during the merging stage, as shown in Table 3.
@@
State represents the current development context, situated within the task and environment. Unlike prior taxonomies (Mozannar et al., 2024; Amoozadeh et al., 2024) that predetermine states or general activities (e.g., “Waiting For Suggestion” or “Prompt Crafting”), we infer states dynamically from a participant’s codebase, observed actions, and tool usage (Research, 2023). We argue that the State dimension serves as the foundation for developer behavior: (1) Developers form an intention based on the current state of the environment. For instance, if the terminal displays a compilation error, the developer may form the intention to resolve the error. (2) The developer then translates this intention into an action, such as editing code or searching for a solution using different supporting tools, while experiencing a specific emotion. (3) This combination of intention, action, and emotion leads a transition to a new state (e.g., implementing a new function or encountering a new bug), creating a feedback loop where the updated state influences the next intention and action.
@@
As developers often switch between subtasks (e.g., writing a new snippet and then running tests to debug), states can overlap as explained by Mozannar et al. (Mozannar et al., 2024). We do not provide a formal state model here; we argue that the state dimension is inherently fluid and context dependent. However, we argue that recognizing the State as a dynamic input is crucial for understanding how developers form intentions and take action within assisted tools and emotional responses during software development.
@@
$\bullet$ Intention reflects a developer’s goal for software development (Yu et al., 2025; AlOmar et al., 2024; Research, 2023). We argue that the Intention is the motivation (e.g., the goal is to do something) behind every one of their physical or digital actions. Hence, Intentions are often invisible, while actions are usually visible. For example, a developer intends to understand the task description when given a task, but the developer’s action could vary from prompting AI tools to understanding or using a search engine to retrieve relevant information on the Internet. Drawing on our pilot study and participant task activity annotations, we identified 12 Intentions as shown in Table 3.
@@
$\bullet$ Action describes how the developer proceeds toward a specific intention. We identified 16 Action categories as shown in Table 3. Some of these (e.g., Writing New Code, Running Tests) are readily visible in screen recordings, while others (e.g., Reading and Comprehension) were partially inferred from user behavior (e.g., scrolling or paused reading) and participant self annotations.
@@
$\bullet$ Supporting tools Indicate whether and how developers use external tools during an action. Prior studies show actions differ significantly when AI is involved. To capture this distinction, we separate Action from Supporting Tools. Given the nature of our study, AI assistants were frequently relied upon to support an Action.
@@
$\bullet$ Emotion is measured by a 7-point valence scale as shown in Table 3. While it is hard to capture instant feelings without external tools such as eye tracking devices or brain cognitive analysis tools, we adopt Russell’s circumplex model of affect (Russell, 1980) to measure emotions in an exploratory view, and we followed the practice of Gardella et al. (Gardella et al., 2024) to adopt 7-level valence scale. We prioritized valence due to its practicality in retrospective annotation (Gardella et al., 2024). For example, a developer might report a negative valence after repeated failed attempts to resolve errors with AI assistants.
@@ 4.2. Validation and Member Checking
To validate our proposed model category set, we employed two validation approaches. First, in the post-survey, we included an optional validation question on a 5-point Likert scale: “The multiple input fields (e.g., three input boxes) capture my thoughts (e.g., intentions and emotions) and actions effectively.” Among the 32 participants who responded, 4 (12%) strongly agreed, 19 (59%) agreed, 5 (16%) neutral, 3 (9%) disagreed, and 1 (3%) strongly disagreed, indicating that most participants considered the input fields reflective of their actual programming behavior. We further conducted member checking (Harper and Cole, 2012; Ford et al., 2017) with the final category set. Among the 20 participants who responded, 15 strongly agreed with our merged labels, while the remaining participants expressed uncertain feelings. Through follow-up inquiries, we found that these 5 participants did not strongly disagree or disagree with the existing categories but rather struggled with the inherent ambiguity of their own latent intentions in a few programming states. They noted that while our labels were the most appropriate available, the “hidden” nature of programming intentions occasionally makes precise self-annotation difficult. This underscores that while S-IASE provides a structured model, capturing the full spectrum of developer programming behavior, especially intricate intention, remains an exploratory challenge.

### 26 Apr 2026 v2

@@ 2.6. Annotation Tool and Procedure
To collect data for the S-IASE model, we developed an annotation tool so participants could divide video recordings into multiple time intervals and annotate each interval based on the model. Hence, we used the implementation of a tool for annotating programming videos developed by Mozannar et al. (Mozannar et al., 2024) as our starting point. We designed the graphical user interface (GUI) and developed features that are essential for our study purpose. For instance, following the pilot study, we improved the user experience of the annotation tool by displaying the options with dropdown menus.
Figure 2 depicts our annotation tool GUI. Once the task execution stage was finished, we asked participants to stop recording their screens. We then provided a demo of the annotation tool. After that, participants annotated their programming session to ensure their recollection of the task remained fresh. Unlike the original tool (Mozannar et al., 2024), ours allows participants to define their own time intervals by setting start and end timestamps.
@@
In the annotation step, we encouraged participants to annotate approximately 20 intervals (each representing a different time range) describing their development states. This guideline was based on the average number of intervals observed during our pilot study. Although we used 20 intervals as a reference for participants, they could freely create as many intervals as they felt necessary. For each time interval, participants needed to annotate all four dimensions of our S-IASE model (described in Section 4). In the initial annotation tool, the predefined categories contained 10 intentions, 14 actions, 8 supporting tools, and 7 emotions labels. They could also add custom labels within each category if they found predefined options insufficient. After data collection, two researchers first reviewed a random subset of 10 screen videos and their corresponding annotated labels to assess the completeness and consistency. After validation, the same researchers conducted a structured merging (Zimmermann, 2016) process to consolidate similar custom categories and refine categories based on participant responses. One researcher reviewed and clustered all custom labels, mapping them onto existing categories or creating new categories as necessary (e.g., adding the intention to set up project environment and the action search). Another researcher independently validated these mappings, and any disagreements (i.e., representing 5% of all categories) were resolved through discussions until consensus was achieved. As shown in Table 3, the newly emerged labels are marked with a (+), indicating that they were added during this merging stage.
@@ 2.7. Pilot Study
Before experimenting on a larger group of participants, we performed a pilot study with eight participants for two purposes: (1) design and refine our developer-AI model derived from our literature review (which was S-IAS at the time, no E), and (2) refine details in our designed user study steps to ensure clarity for participants. The pilot study became pivotal for answering RQ1 because we discovered through piloting that, in addition to intention and action, emotion emerged as a paramount dimension participants repeatedly alluded to. For example, “I am so frustrated that ChatGPT is granting permissions to all the repos via the REST API as opposed to having a more privacy-centric view in its answers.” It became clear to us from the pilots that participants demonstrated a fluctuating level of emotions while using AI tools, which is important to consider if we want to improve and strengthen prior characterizations of developer-AI interaction during programming. This finding is in line with prior work that has studied the impact of emotion on developer productivity (Madampe et al., 2025; Cheng et al., 2023; Graziotin et al., 2015; Girardi et al., 2020).
As a result, we made a major refinement to our model, adding emotion as a dimension. This resulted in four dimensions that formed the foundation for our S-IASE model. Additionally, piloting helped identify annotation categories for each of these four dimensions. For example, the intention To Clarify Requirements/Ambiguities emerged when participants expressed uncertainty regarding task instructions. Actions such as Logging and tools like Database Tools were added to better capture task-specific activities observed during the piloting sessions.
Piloting also allowed us to refine our task instructions to improve clarity and feasibility. For example, the 25-minute task duration was determined based on the performance of pilot participants, who demonstrated that the core requirements of both PyT and JavaT could be addressed within this time-boxed window. For JavaT, we allowed participants to choose between Maven and standard package imports to establish a JDBC connection, reducing setup complexity. We also improved the annotation tool UI by making it into a single page application with all the annotation categories for a given time interval in a scrollable panel. To prevent data contamination, we confirmed that pilot participants were not included in the main study results.
@@ 2.8. Interview
As the last stage of our user study, we invited participants to our follow-up interview. Before the interview, we proposed a set of general interview questions from multiple perspectives.
@@
For constructing our interview scripts, we mainly focused on what they thought, why they behaved in certain ways, how they felt during the experiments, and what they had learned. Therefore, the interview scripts were constructed in three layers. First, we designed general questions on developer–AI interaction inspired by prior studies (e.g., (Mozannar et al., 2024; Barke et al., 2023; Li et al., 2024; Liang et al., 2024; Choudhuri et al., 2025)) and adopted Bloom’s taxonomy (Bloom et al., 1956; Anderson and Krathwohl, 2001) to probe what participants felt they had learned during the task. Second, we focused on cognitive biases and cognitive load under time pressure, using responses from validated instruments such as the NASA-TLX (Hart, 2006) and self-efficacy scales as anchors for follow-up questions (e.g., “You rated mental demand as 5, what was happening in that moment?”). Third, we included questions on the usability of our annotation tool, asking participants to evaluate it and whether it accurately captured their intentions, actions, and emotions. Then, for each interview, we carefully watched the participant’s uploaded video and read their response history to tailor participant-specific follow-up “why” questions, such as “Why did you switch from using an AI tool to searching Stack Overflow at this point?”
@@
Among the 76 participants, 14 (18%) volunteered to be interviewed. Each interview took around 30 minutes with two authors and one participant. Our interviews followed a semi-structured format where we used our prepared interview questions as a general guideline, but allowed flexibility to explore topics that emerged during an interview. After obtaining permission from our interviewees, interview audio was recorded and transcribed by Otter.ai (50). To ensure that no errors were made in the transcription, one author manually verified the transcript for each interview.
@@
The same author who conducted each interview and transcription afterwards performed basic data analysis from socio-technical grounded theory (STGT) (Hoda, 2024). We chose STGT for data analysis due to its grounding in the socio-technical context. Part of the reason STGT was developed was due to “the core SE practice of programming is not strictly technical, rather, it is socio-technical, involving intensive human-human collaboration and coordination and human-technology interactions.” (Hoda, 2021) With the addition of AI assistance to programming, the subject matter we studied is fittingly in the socio-technical domain. In contrast to traditional grounded theory methods, which are dedicated to theory development and do not support limited applications for data analysis, STGT can support data analysis to generate rich descriptive themes and categories. These themes and categories helped supplement our proposed developer-AI model and findings about observed developer behavior. Furthermore, we adopted a reflexive, researcher-led approach to conduct STGT from an interpretivist perspective (Hoda, 2024). The first author served as the primary analytic instrument, conducting open coding, constant comparison, and iterative memoing across all interview transcripts. This process resulted in a range of themes. For this paper, we present themes most aligned with our research focus, such as “Trust but Verify.” Member checking with a subset of 6 interviewees helped confirm the resonance and validity of the presented interpretations.
@@ 3. Task Completion Results
Completed Correct Completed Correct Completed Correct Completed Correct
@@
Completed Correct Completed Correct Completed Correct Completed Correct Completed Correct
@@
Table 1 details how our participants performed on each of the subtasks for PyT and JavaT. We distinguished between Completed (the participant performed the required actions and finished the specific subtask) and Correct (the outcome of the subtask was verified to be functionally accurate). A participant achieved “All Subtasks Correct” if they completed and correctly solved all subtasks within 25 minutes. Among 39 participants for PyT, only 8 were successful in solving all subtasks, while only 2 out of 37 were successful for JavaT. All 10 participants who correctly solved all their assigned subtasks were from AI-assisted groups. We found that 42% (32/76) of all participants successfully reached the final subtask of their assigned activity (56% for PyT and 27% for JavaT). However, the overall rate for full functional correctness remained at 13% (10/76). This suggests that while the tasks were possible to navigate within the time limit, the complexity of API and database interactions presented a substantial challenge to functional accuracy. The lower success rate in the Java task, compared to the Python task, further indicates that environment setup and multi-step database operations imposed a heavier correctness burden on developers. Moreover, we found that the AI group achieved a higher completion rate than the non-AI group for each subtask. The results also indicate that while AI markedly boosted both completion and correctness, the complexity of the tasks presented a substantial challenge within the 25-minute limit.
On average, participants in AI-assisted groups adopted more than one AI assistant (1.37) during the experiment. ChatGPT was by far the most popular, used by 38 participants, with an additional 13 participants relying on the paid version. 12 participants used GitHub Copilot as their AI assistant in IDE. This distribution indicates multiple assistants’ strategies for adopting AI for task completion.
@@ 4.1. Outcome
Our participants were able to add labels to the model as needed. After data collection and structured merging, we indentified 12 intentions, 16 actions, 8 supporting tools, and 7 emotions. Newly emerged labels are marked with (+) to indicate they were added during the merging stage, as shown in Table 3.
@@
State represents the current development context, situated within the task and environment. Unlike prior taxonomies (Mozannar et al., 2024; Amoozadeh et al., 2024) that predetermine states or general activities (e.g., “Waiting For Suggestion” or “Prompt Crafting”), we infer states dynamically from a participant’s codebase, observed actions, and tool usage (Research, 2023). We argue that the State dimension serves as the foundation for developer behavior: (1) Developers form an intention based on the current state of the environment. For instance, if the terminal displays a compilation error, the developer may form the intention to resolve the error. (2) The developer then translates this intention into an action, such as editing code or searching for a solution using different supporting tools, while experiencing a specific emotion. (3) This combination of intention, action, and emotion leads to a transition to a new state (e.g., implementing a new function or encountering a new bug), creating a feedback loop where the updated state influences the next intention and action.
@@
As developers often switch between subtasks (e.g., writing a new snippet and then running tests to debug), states can overlap (Mozannar et al., 2024). We do not provide a formal state model here; we argue that the state dimension is inherently fluid and context-dependent. However, recognizing the State as a dynamic input is crucial for understanding how developers form intentions, take action within assisted tools, and experience emotional responses during software development.
@@
$\bullet$ Intention reflects a developer’s goal for software development (Yu et al., 2025; AlOmar et al., 2024; Research, 2023). We argue that the Intention is the motivation (e.g., the goal is to do something) behind every one of their physical or digital actions. Hence, Intentions are often invisible, while actions are usually visible. For example, a developer intends to understand the task description when given a task, but the developer’s action could vary from prompting AI tools to understanding or using a search engine to retrieve relevant information on the Internet. Drawing on our pilot study and participant task activity annotations, we identified 12 Intentions, as shown in Table 3.
@@
$\bullet$ Action describes how a developer proceeds toward a specific intention. We identified 16 Action categories, as shown in Table 3. Some of these (e.g., Writing New Code, Running Tests) are readily visible in screen recordings, while others (e.g., Reading and Comprehension) are partially inferred from user behavior (e.g., scrolling or paused reading) and participant self-annotations.
@@
$\bullet$ Supporting tools indicate whether and how a developer uses external tools during an action. Prior studies show that actions differ significantly when AI is involved. To capture this distinction, we separate Action from Supporting Tools. Given the nature of our study, AI assistants were frequently relied upon to support an Action.
@@
$\bullet$ Emotion is measured by a 7-point valence scale as shown in Table 3. While it is hard to capture instant feelings without external tools such as eye-tracking devices or brain cognitive analysis tools, we adopted Russell’s circumplex model of affect (Russell, 1980) to measure emotions from an exploratory perspective, and we followed the practice of Gardella et al. (Gardella et al., 2024) to adopt a 7-point valence scale. We prioritized valence due to its practicality in retrospective annotation (Gardella et al., 2024). For example, a developer might report a negative valence after repeated failed attempts to resolve errors with AI assistants.
@@ 4.2. Validation and Member Checking
To validate our proposed model category set, we employed two validation approaches. First, in the post-survey, we included an optional validation question on a 5-point Likert scale: “The multiple input fields (e.g., three input boxes) capture my thoughts (e.g., intentions and emotions) and actions effectively.” Among the 32 participants who responded, 4 (12%) strongly agreed, 19 (59%) agreed, 5 (16%) were neutral, 3 (9%) disagreed, and 1 (3%) strongly disagreed, indicating that most participants considered the input fields reflective of their actual programming behavior. We further conducted member checking (Harper and Cole, 2012; Ford et al., 2017) with the final category set. Among the 20 participants who responded, 15 strongly agreed with our merged labels, while the remaining participants expressed uncertainty. Through follow-up inquiries, we found that these 5 participants did not strongly disagree or disagree with the existing categories but rather struggled with the inherent ambiguity of their own latent intentions in a few programming states. They noted that while our labels were the most appropriate available, the “hidden” nature of programming intentions occasionally made precise self-annotation difficult. This underscores that while S-IASE provides a structured model, capturing the full spectrum of developer programming behavior, especially intricate intentions, remains an exploratory challenge.